# Revisiting Non-Acyclic GFlowNets in Discrete Environments

### v1

@@ 1 Introduction
Generative Flow Networks (GFlowNets, Bengio et al., 2021) are models that aim to sample discrete objects from distributions known up to proportionality constant. They operate by constructing an object through a sequence of stochastic transitions defined by a forward policy. GFlowNets have been successfully applied in various areas, starting from molecule generation (Bengio et al., 2021) and biological sequence design (Jain et al., 2022) to combinatorial optimization (Zhang et al., 2023) and fine-tuning of diffusion models (Uehara et al., 2024). The detailed theoretical foundations of GFlowNets in discrete environments were developed in (Bengio et al., 2023). While the majority of GFlowNet literature consider the discrete setting, it is possible to apply the methodology of continuous GFlowNets (Lahlou et al., 2023) to sampling problems on more general spaces.
@@
The main idea behind the generation process in GFlowNets lies in sampling trajectories in the appropriately constructed directed acyclic graph environment instead of working directly in the object space. A standard intuition behind this process is a sequence of actions applied in order to construct a composite object from ”blocks”. One of the limitations of this setting is that it requires acyclicity. While this limitation can be naturally interpreted in, e.g., molecular generation setting, it can potentially restrict the applicability of GFlowNets in other scenarios, e.g. in standard RL environments or for sampling objects with intrinsic symmetries which are better modeled in a non-acyclic environment, like a Cayley graph of Rubick’s cube. To the best of our knowledge, methodological aspects of working with non-acyclic graphs in GFlowNets were previously considered only in the recent work (Brunswic et al., 2024). The latter paper, similarly to (Lahlou et al., 2023), uses the machinery of measurable spaces and measure theory, which is harder to build new extensions and methodology upon. We believe that simplicity is a key merit of the theory behind discrete GFlowNets (Bengio et al., 2023) when compared to their general state counterparts. Thus, the main aim of our paper is to revisit the theory of non-acyclic GFlowNets with a discrete state space, simplifying the constructions of (Brunswic et al., 2024) and providing additional theoretical and methodological insights into training GFlowNets in this setting. The main contributions of the paper can be summarized as follows:
@@ 2.1 GFlowNets
This section presents necessary notations and theoretical background on GFlowNets. GFlowNets treat the problem of sampling from a probability distribution over discrete space $\mathcal{X}$ as a sequential decision-making process in a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{E})$, where $\mathcal{S}$ is a state space and $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ is a set of edges (or transitions). There is a special initial state $s_{0}$ with no incoming edges and a special sink state $s_{f}$ with no outgoing edges. The commonly used variant of notation does not include a sink state $s_{f}$, yet we prefer to use a variant with $s_{f}$, since it was also used in the previous work on non-acyclic GFlowNets (Brunswic et al., 2024) and leads to a more intuitive construction. Let $\mathcal{T}$ be a set of all trajectories $\tau=\mathopen{}\mathclose{{}\left(s_{0}\to s_{1}\to\ldots\to s_{n_{\tau}}\to s_{f}}\right)$ from $s_{0}$ to $s_{f}$, where we use $n_{\tau}$ to denote the length of the trajectory $\tau$. We use a convention $s_{n_{\tau}+1}=s_{f}$. We say that $\tau$ terminates in $s$ if its last transition is $s\to s_{f}$. Such transitions are called terminating, and the states that have an outgoing edge into $s_{f}$ are called terminal states. The set of terminal states coincides with $\mathcal{X}$, and the probability distribution of interest $\mathcal{R}(x)/\mathcal{Z}$ is defined on it, where $\mathcal{R}(x)$ is called GFlowNet reward and $\mathcal{Z}=\sum_{x\in\mathcal{X}}\mathcal{R}(x)$ is an unknown normalizing constant. In addition, for any state $s$, we denote $\operatorname{in}(s)$ to be the set of states $s^{\prime}$ such that there is an edge $s^{\prime}\to s\in\mathcal{E}$ (parents), and $\operatorname{out}(s)$ to be the set of states $s^{\prime}$ such that there is an edge $s\to s^{\prime}\in\mathcal{E}$ (children).
@@
All of these objects are completely and uniquely specified if one fixes either 1) edge flow $\mathcal{F}(s\to s^{\prime})$, 2) initial flow $\mathcal{F}(s_{0})$ and $\mathcal{P}_{\mathrm{F}}$, 3) initial flow $\mathcal{F}(s_{0})$ and $\mathcal{P}_{\mathrm{B}}$. If flows satisfy $\mathcal{F}(s\to s_{f})=\mathcal{R}(s)$, trajectory distribution defined by the corresponding forward policy will satisfy reward matching condition (Bengio et al., 2023). In practice, a neural network is used to parameterize the forward policy (and, optionally, the backward policy and the flows). Then, it is trained to minimize some loss function that would enforce the reward matching condition. For example, Detailed Balance loss (Bengio et al., 2023) is defined on all transitions $s\to s^{\prime}\in\mathcal{E}$ as:
@@ 2.2 GFlowNets in Non-Acyclic Environments
Brunswic et al. (2024) state that fundamental results of GFlowNet theory also apply in the case when the environment graph $\mathcal{G}$ may contain cycles, and all definitions from the acyclic case remain valid and extend to the non-acyclic case. However, we will further show that this is not exactly true, e.g., flows cannot be consistently defined as unnormalized visitation probabilities.
More specifically, Brunswic et al. (2024) argue that if (3) holds for an edge flow, as well as $\mathcal{F}(s\to s_{f})=\mathcal{R}(s)$ for terminating transitions, the forward policy induced by the flow $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\frac{\mathcal{F}(s\to s^{\prime})}{\mathcal{F}(s)}$ satisfies reward matching condition. Thus, standard GFlowNet losses, such as Flow Matching (FM, Bengio et al., 2021), Detailed Balance (DB, Bengio et al., 2023), and Trajectory Balance (TB, Malkin et al., 2022) can be applied in non-acyclic environments.
However, Brunswic et al. (2024) point out that the main distinction between non-acyclic and acyclic GFlowNets is that in the non-acyclic setting, expected trajectory length $\mathbb{E}[n_{\tau}]$ (denoted as a sampling time in (Brunswic et al., 2024)) can be arbitrary large because of the cycles, while in the acyclic setting it is always bounded. To tackle this issue, a concept of loss stability is introduced. A loss is called stable if adding a constant to the flow along a cycle can never decrease the loss, and otherwise, it is called unstable (Definition 3). It is shown that FM, DB, and TB are unstable (Theorem 3), which can lead to arbitrarily large sampling time when utilized for training. In contrast, a family of losses that are provably stable is presented (Theorem 4). Moreover, the authors show that there are stable variants of FM and DB losses, such as stable DB loss, which we denote as SDB
$\begin{split}\mathcal{L}_{\texttt{SDB}}(s\to s^{\prime})&\triangleq\log(1+\varepsilon\Delta^{2}(s,s^{\prime},\theta))(1+\eta\mathcal{F}_{\theta}(s))\,,\\ \ \Delta(s,s^{\prime},\theta)&\triangleq\mathcal{F}_{\theta}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}|s,\theta)-\mathcal{F}_{\theta}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s|s^{\prime},\theta)\,,\end{split}$ (5)
@@ 3.1 Environment
All definitions regarding the environment can be introduced similarly to the setting of acyclic GFlowNets (see Section 2.1) with one main difference: graph $\mathcal{G}$ can now contain cycles. We make two technical assumptions on the structure of $\mathcal{G}$:
@@
1) There is a special initial state $s_{0}$ with no incoming edges and a special sink state $s_{f}$ with no outgoing edges; 2) For any state $s\in\mathcal{S}$ there exists a path from $s_{0}$ to $s$ and a path from $s$ to $s_{f}$.
@@
A sequence $\tau=(s_{0}\to s_{1}\to\ldots\to s_{n_{\tau}}\to s_{n_{\tau}+1}=s_{f})$ is called a trajectory of length $n_{\tau}\in\mathbb{N}$ if all transitions $s_{t}\to s_{t+1}\in\mathcal{E},t\in\{0,\ldots,n_{\tau}\}$. Then $\mathcal{T}$ is a set of all finite length trajectories which start in $s_{0}$ and finish in $s_{f}$.
In the above definition and further, we use a convention $s_{n_{\tau}+1}=s_{f}$. The main difference with acyclic GFlowNets is that $\mathcal{T}$ can be infinite, and $\mathcal{T}$ can contain trajectories of arbitrary length.
@@ 3.2 Backward Policy and Trajectory Distribution
Let $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ be a backward policy, such that $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for any edge $(s\to s^{\prime})\in\mathcal{E}$. Then
@@
In fact, the condition $\mathcal{P}_{\mathrm{B}}(s^{\prime}\mid s)>0$ together with 3.1 allows us to ensure that the sequence $s_{i}$ is a finite state-space absorbing Markov chain. Given this assumption, the proof of Lemma 3.4 almost coincides with the proof of the fact that the states of such a Markov chain are non-recurrent, see, e.g., (Levin & Peres, 2017). For completeness, we provide the proof in Appendix A.1.
@@ 3.3 State and Edge Flows
The number above each edge $(s\to s^{\prime})$ is $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Consider the distribution $\mathcal{P}(\tau)$ defined by (7). Let us plot the visitation probability for each edge:
@@
It is easy to check that the flow matching constraints (3) are now satisfied. Next, we formally define:
@@
Let $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ be a backward policy, such that $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for any edge $(s\to s^{\prime})\in\mathcal{E}$. Then, given a final flow $\mathcal{F}(s_{f})>0$, we define state and edge flows as
@@
    $\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ for any $(s\to s^{\prime})\in\mathcal{E}$.
@@
Let $\mathcal{F}:\mathcal{E}\to\mathbb{R}_{>0}$ be a function that satisfies flow matching conditions (3). Define the corresponding backward policy by the relation
@@ 3.4 Forward Policy and Detailed Balance
It is well-known in acyclic GFlowNets theory (Bengio et al., 2023) that there exists a unique forward policy $\mathcal{P}_{\mathrm{F}}$ for any backward policy $\mathcal{P}_{\mathrm{B}}$ that induces the same probability distribution over $\mathcal{T}$. The main implication of this fact is that by fixing rewards $\mathcal{R}(x),x\in\mathcal{X}$ and a backward policy $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ for each state $s^{\prime}\in\mathcal{S}\setminus\{s_{0},s_{f}\}$, one automatically fixes a trajectory distribution $\mathcal{P}(\tau)$ that satisfies the reward matching condition (Malkin et al., 2022). However, sampling from such distribution is intractable since it requires starting from a terminal state sampled from the reward distribution. Thus, during GFlowNet training, one tries to find a forward policy, which always allows tractable sampling of trajectories by construction, that will match this trajectory distribution $\mathcal{P}(\tau)$. One can note that this bears similarities with hierarchical variational inference (Malkin et al., 2023b). In the following proposition, we show that this result holds also for non-acyclic GFlowNets.
@@
The proof is provided in Appendix A.4. Conversely, the following proposition shows that if a triplet $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$, $\mathcal{P}_{\mathrm{B}}$ satisfies detailed balance conditions (2), it will be consistent with all previous definitions and propositions.
@@
Let $\mathcal{F}\colon\mathcal{S}\to\mathbb{R}_{>0}$, and let $\mathcal{P}_{\mathrm{F}}(s^{\prime}|s)>0$, $\mathcal{P}_{\mathrm{B}}(s|s^{\prime})>0$ be forward and backward policies, such that detailed balance conditions (2) are satisfied. Then $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$ induce the same trajectory distribution:

### v2

@@ 1 Introduction
Generative Flow Networks (GFlowNets, Bengio et al., 2021) are models that aim to sample discrete objects from distributions known proportionally up to a constant. They operate by constructing an object through a sequence of stochastic transitions defined by a forward policy. GFlowNets have been successfully applied in various areas, starting from molecule generation (Bengio et al., 2021; Shen et al., 2024; Koziarski et al., 2024; Cretu et al., 2025) and biological sequence design (Jain et al., 2022; Kim et al., 2024) to combinatorial optimization (Zhang et al., 2023a, b; Kim et al., 2025) and fine-tuning of large language models and diffusion models (Hu et al., 2023; Venkatraman et al., 2024; Zhang et al., 2025; Uehara et al., 2024; Lee et al., 2025). The detailed theoretical foundations of GFlowNets in discrete environments were developed in (Bengio et al., 2023). While the majority of GFlowNet literature considers the discrete setting, it is possible to apply the methodology of continuous GFlowNets (Lahlou et al., 2023) to sampling problems on more general spaces.
The main idea behind the generation process in GFlowNets lies in sampling trajectories in the appropriately constructed directed acyclic graph environment instead of working directly in the object space. A standard intuition behind this process is a sequence of actions applied in order to construct a composite object from ”blocks”. One of the limitations of this setting is that it requires acyclicity. While this limitation can be naturally interpreted in, e.g., molecule generation setting, it can confine the practical design of GFlowNet environments, as well as restrict the applicability of GFlowNets in other scenarios. A motivational example for non-acyclic GFlowNets presented by (Brunswic et al., 2024) is related to modeling distributions over objects with intrinsic symmetries. Consider a class of environments where states are elements of some group, e.g. symmetric group or Rubik’s Cube group. The transitions are given via a generating set of this group, thus corresponding to applying the group operation on the current state and some element of the generating set, which leads to existence of cycles. While in some cases an acyclic environment can be designed to generate group elements, such environments of “algebraic” origin naturally contain cycles, thus falling under the area of our study. In addition, there is a growing body of work which connects GFlowNets and Reinforcement Learning (Tiapkin et al., 2024; Mohammadpour et al., 2024; Deleu et al., 2024; He et al., 2024). Most RL environments contain cycles, thus understanding how GFlowNets can be applied in cyclic environments and connecting them to RL in such case can be a crucial step towards further bridging two research fields.
@@
To the best of our knowledge, methodological aspects of working with non-acyclic environments in GFlowNets were previously considered only in the recent work of (Brunswic et al., 2024). The latter paper, similarly to (Lahlou et al., 2023), uses the machinery of measurable spaces and measure theory, which is harder to build new extensions and methodology upon. We believe that simplicity is a key merit of the theory behind discrete GFlowNets (Bengio et al., 2023) when compared to their general state counterparts. Thus, the main aim of our paper is to revisit the theory of non-acyclic GFlowNets within a discrete state space, simplifying the constructions of (Brunswic et al., 2024) and providing additional theoretical and methodological insights into training GFlowNets in this setting. The main contributions of the paper can be summarized as follows:
@@ 2.1 GFlowNets
This section presents necessary notations and theoretical background on GFlowNets. GFlowNets treat the problem of sampling from a probability distribution over discrete space $\mathcal{X}$ as a sequential decision-making process in a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{E})$, where $\mathcal{S}$ is a finite state space and $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{S}$ is a finite set of edges (or transitions). There is a special initial state $s_{0}$ with no incoming edges and a special sink state $s_{f}$ with no outgoing edges. The commonly used variant of notation does not include a sink state $s_{f}$, yet we prefer to use a variant with $s_{f}$, since it was also used in the previous work on non-acyclic GFlowNets (Brunswic et al., 2024) and leads to a more intuitive construction. Let $\mathcal{T}$ be a set of all trajectories $\tau=\mathopen{}\mathclose{{}\left(s_{0}\to s_{1}\to\ldots\to s_{n_{\tau}}\to s_{f}}\right)$ from $s_{0}$ to $s_{f}$, where we use $n_{\tau}$ to denote the length of the trajectory $\tau$. We use a convention $s_{n_{\tau}+1}=s_{f}$. We say that $\tau$ terminates in $s$ if its last transition is $s\to s_{f}$. Such transitions are called terminating, and the states that have an outgoing edge into $s_{f}$ are called terminal states. The set of terminal states coincides with $\mathcal{X}$, and the probability distribution of interest $\mathcal{R}(x)/\mathcal{Z}$ is defined on it, where $\mathcal{R}(x)>0$ is called GFlowNet reward and $\mathcal{Z}=\sum_{x\in\mathcal{X}}\mathcal{R}(x)$ is an unknown normalizing constant. In addition, for any state $s$, we denote $\operatorname{in}(s)$ to be the set of states $s^{\prime}$ such that there is an edge $s^{\prime}\to s\in\mathcal{E}$ (parents), and $\operatorname{out}(s)$ to be the set of states $s^{\prime}$ such that there is an edge $s\to s^{\prime}\in\mathcal{E}$ (children).
@@
All of these objects are completely and uniquely specified if one fixes either 1) edge flow $\mathcal{F}(s\to s^{\prime})$, 2) initial flow $\mathcal{F}(s_{0})$ and $\mathcal{P}_{\mathrm{F}}$, 3) initial flow $\mathcal{F}(s_{0})$ and $\mathcal{P}_{\mathrm{B}}$. If flows satisfy $\mathcal{F}(s\to s_{f})=\mathcal{R}(s)$, trajectory distribution defined by the corresponding forward policy will satisfy the reward matching condition (Bengio et al., 2023). In practice, a neural network is used to parameterize the forward policy (and, optionally, the backward policy and the flows). Then, it is trained to minimize some loss function that would enforce the reward matching condition. For example, Detailed Balance loss (Bengio et al., 2023) is defined on all transitions $s\to s^{\prime}\in\mathcal{E}$ as:
@@ 2.2 GFlowNets in Non-Acyclic Environments
(Brunswic et al., 2024) state that fundamental results of GFlowNet theory also apply in the case when the environment graph $\mathcal{G}$ may contain cycles, and all definitions from the acyclic case remain valid and extend to the non-acyclic case. However, we will further show that this is not exactly true, e.g., flows cannot be consistently defined as unnormalized visitation probabilities.
More specifically, (Brunswic et al., 2024) argue that if (3) holds for an edge flow, as well as $\mathcal{F}(s\to s_{f})=\mathcal{R}(s)$ for terminating transitions, the forward policy induced by the flow $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\frac{\mathcal{F}(s\to s^{\prime})}{\mathcal{F}(s)}$ satisfies the reward matching condition. Thus, standard GFlowNet losses, such as Flow Matching (FM, Bengio et al., 2021), Detailed Balance (DB, Bengio et al., 2023), and Trajectory Balance (TB, Malkin et al., 2022) can be applied in non-acyclic environments.
However, (Brunswic et al., 2024) point out that the main distinction between non-acyclic and acyclic GFlowNets is that in the non-acyclic setting, expected trajectory length $\mathbb{E}[n_{\tau}]$ (denoted as a sampling time in (Brunswic et al., 2024)) can be arbitrarily large because of the cycles, while in the acyclic setting it is always bounded. To tackle this issue, a concept of loss stability is introduced. A loss is called stable if adding a constant to the flow along a cycle can never decrease the loss, and otherwise, it is called unstable (Definition 3). It is shown that FM, DB, and TB are unstable (Theorem 3), which can lead to arbitrarily large sampling time when utilized for training. In contrast, a family of losses that are provably stable is presented (Theorem 4). Moreover, the authors show that there are stable variants of FM and DB losses, such as stable DB loss, which we denote as SDB:
$\mathcal{L}_{\texttt{SDB}}(s\to s^{\prime})&\triangleq\log(1+\varepsilon\Delta^{2}(s,s^{\prime},\theta))(1+\eta\mathcal{F}_{\theta}(s))\,,\\ \ \Delta(s,s^{\prime},\theta)&\triangleq\mathcal{F}_{\theta}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}|s,\theta)-\mathcal{F}_{\theta}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s|s^{\prime},\theta)\,,\end{aligned}$
@@ 3.1 Environment
All definitions regarding the environment can be introduced similarly to the setting of acyclic GFlowNets (see Section 2.1) with one main difference: graph $\mathcal{G}$ can now contain cycles. In addition to finiteness, we make two technical assumptions on the structure of $\mathcal{G}$:
@@
0) graph $\mathcal{G}$ is finite; 1) There is a special initial state $s_{0}$ with no incoming edges and a special sink state $s_{f}$ with no outgoing edges; 2) For any state $s\in\mathcal{S}$ there exists a path from $s_{0}$ to $s$ and a path from $s$ to $s_{f}$.
@@
A sequence $\tau=(s_{0}\to s_{1}\to\ldots\to s_{n_{\tau}}\to s_{n_{\tau}+1}=s_{f})$ is called a trajectory of length $n_{\tau}\in\mathbb{N}$ if all transitions $s_{t}\to s_{t+1}\in\mathcal{E},t\in\{0,\ldots,n_{\tau}\}$. Then $\mathcal{T}$ is a set of all finite-length trajectories that start in $s_{0}$ and finish in $s_{f}$.
In the above definition and further, we use a convention $s_{n_{\tau}+1}=s_{f}$. While $\mathcal{G}$ itself is always finite, the main difference with acyclic GFlowNets is that $\mathcal{T}$ can be infinite, and $\mathcal{T}$ can contain trajectories of arbitrary length.
@@ 3.2 Backward Policy and Trajectory Distribution
Let $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ be a backward policy, such that $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for any edge $s\to s^{\prime}\in\mathcal{E}$. Then
@@
In fact, the condition $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ together with 3.1 allows us to ensure that the sequence $s_{i}$ is a finite state-space absorbing Markov chain. Given this assumption, the proof of Lemma 3.4 almost coincides with the proof of the fact that the states of such a Markov chain are transient, see, e.g., (Douc et al., 2018). For completeness, we provide the proof in Appendix A.1.
@@ 3.3 State and Edge Flows
The number above each edge $s\to s^{\prime}$ is $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$. Consider the distribution $\mathcal{P}(\tau)$ defined by (7). Let us plot the visitation probability for each edge:
@@
It is easy to check that the flow matching conditions (3) are now satisfied. Next, we formally define:
@@
Let $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ be a backward policy, such that $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ for any edge $s\to s^{\prime}\in\mathcal{E}$. Then, given a final flow $\mathcal{F}(s_{f})>0$, we define state and edge flows as
@@
    $\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ for any $s\to s^{\prime}\in\mathcal{E}$.
@@
Let $\mathcal{F}:\mathcal{E}\to\mathbb{R}_{>0}$ be a function that satisfies the flow matching conditions (3). Define the corresponding backward policy by the relation
@@ 3.4 Forward Policy and Detailed Balance
It is well-known in acyclic GFlowNets theory (Bengio et al., 2023) that there exists a unique forward policy $\mathcal{P}_{\mathrm{F}}$ for any backward policy $\mathcal{P}_{\mathrm{B}}$ that induces the same probability distribution over $\mathcal{T}$. The main implication of this fact is that by fixing rewards $\mathcal{R}(x),x\in\mathcal{X}$ and a backward policy $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ for each state $s^{\prime}\in\mathcal{S}\setminus\{s_{0},s_{f}\}$, one automatically fixes a trajectory distribution $\mathcal{P}(\tau)$ that satisfies the reward matching condition (Malkin et al., 2022). However, sampling from a such distribution is intractable since it requires starting from a terminal state sampled from the reward distribution. Thus, during GFlowNet training, one tries to find a forward policy, which always allows tractable sampling of trajectories by construction, that will match this trajectory distribution $\mathcal{P}(\tau)$. One can note that this bears similarities with hierarchical variational inference (Malkin et al., 2023). In the following proposition, we show that this result also holds for non-acyclic GFlowNets.
@@
The proof is provided in Appendix A.4. Conversely, the following proposition shows that if a triplet $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$, $\mathcal{P}_{\mathrm{B}}$ satisfies the detailed balance conditions (2), it will be consistent with all previous definitions and propositions.
@@
Let $\mathcal{F}\colon\mathcal{S}\to\mathbb{R}_{>0}$, and let $\mathcal{P}_{\mathrm{F}}(s^{\prime}|s)>0$, $\mathcal{P}_{\mathrm{B}}(s|s^{\prime})>0$ be forward and backward policies, such that the detailed balance conditions (2) are satisfied. Then $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$ induce the same trajectory distribution: